OpenStudy (anonymous):

solve the equation: 2sin^2(x)cos(x)=cos(x) on the interval 0,2pie

OpenStudy (anonymous):

it is the only other question i need answers to

OpenStudy (jdoe0001):

solve for \(\bf sin^2(x)\) what would that give you?

OpenStudy (anonymous):

hint maybe

OpenStudy (jdoe0001):

hint: linear simplification

OpenStudy (anonymous):

i am going into the 12th grade and i dont get what u r saying do you think you can help me a little bit more please

OpenStudy (jdoe0001):

think about of sin(x) = a and cos(x) = b so what you really have is -> \(\bf 2a^2b=b\)

OpenStudy (jdoe0001):

if you were to solve for \(\bf a^2\) what would that give?

OpenStudy (anonymous):

umh...0/2 maybe

OpenStudy (jdoe0001):

hmmm \(\bf \cfrac{cos(x)}{cos(x)}=0?\)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

am i wrong

OpenStudy (jdoe0001):

so \(\bf 2sin^2(x)cos(x)=cos(x)\implies 2sin^2(x)=\cfrac{\cancel{ cos(x) }}{\cancel{ cos(x) }}\implies sin^2(x)=\cfrac{1}{2} \\ \quad \\ \sqrt{sin^2(x)}=\pm\sqrt{\cfrac{1}{2}}\implies sin(x)=\pm\sqrt{\cfrac{1}{2}} \\ \quad \\ \textit{taking }sin^{-1}\qquad thus \\ \quad \\ sin^{-1}[sin(x)]=sin^{-1}\left(\pm\sqrt{\cfrac{1}{2}}\right)\implies ?\)

OpenStudy (jdoe0001):

nope.. 1 is correct... so it wouldn't be 0/2, it'd be 1/2

OpenStudy (jdoe0001):

one thing to keep in mind I'd say. is that \(\bf sin^2(x)\to [sin(x)]^2\quad thus \quad \sqrt{sin^2(x)}\implies \sqrt{[sin(x)]^2}\to sin(x)\)

OpenStudy (anonymous):

that is....

OpenStudy (jdoe0001):

so.... \(\bf sin^{-1}[sin(x)]=sin^{-1}\left(\pm\sqrt{\cfrac{1}{2}}\right)\implies ?\)

OpenStudy (anonymous):

can u maybe give me the answer

OpenStudy (jdoe0001):

well.... skipping material is simple going to create greater issues so..... "help" here is defined as help you understand it

OpenStudy (anonymous):

ok wouldnt the answer be like um a degree

OpenStudy (jdoe0001):

yes... it does.. but you first need to solve for "x"

OpenStudy (anonymous):

put it all to x or what

OpenStudy (jdoe0001):

well..... say for example... what's the \(\bf sin(90^o)?\)

OpenStudy (anonymous):

\[\sqrt{4}/2\]

OpenStudy (jdoe0001):

hmm welll yes.. and say now what would be \(\bf sin^{-1}\left(\frac{\sqrt{4}}{2}\right)?\)

OpenStudy (anonymous):

90 degrees

OpenStudy (anonymous):

or pie/2

OpenStudy (anonymous):

thanx i think i got it

OpenStudy (jdoe0001):

so... you see.... then that means that \(\bf sin(90^o)={\color{brown}{ \frac{\sqrt{4}}{2}}}\qquad sin^{-1}\left({\color{brown}{ \frac{\sqrt{4}}{2}}}\right)\implies sin^{-1}\left(sin(90^o)\right)\iff 90^o\)